# RECAST: Reparameterized, Compact weight Adaptation for Sequential Tasks

### v1

@@ 2.3 Parametric Scaling Strategy
RECAST’s design is governed by three parameters: the number of groups ($G$), templates per group ($n$), and coefficient sets ($K$), balancing model expressiveness and efficiency. Increasing $G$ enhances specialization, while larger $n$ boosts expressiveness within groups. The parameter-sharing scheme allows $K$ to expand weight configurations without significantly increasing parameters. This design achieves substantial memory savings, approximated as: $\mathbb{S}=L\times M_{l}\times d^{2}-(G\times n\times d^{2}+L\times M_{l}\times n\times K)$. Here, $M_{l}$ refers to the number of target modules per layer, and we can approximate the savings as: $\mathbb{S}\approx L\times M_{l}\times d^{2}-G\times n\times d^{2}=d^{2}(L\times M_{l}-G\times n)$The savings primarily stem from replacing $L\times M_{l}$ weight matrices with $G\times n$ templates. For experiments, $G$ and $n$ are calibrated to match the baseline model’s parameter count but can be adjusted for desired compression. Additionally, multiple coefficient sets enable a combinatorially large number of weight matrices ($n^{K}$) per module, covering extensive parameter spaces even with small $n$. The total learnable coefficients are $\sum_{l=1}^{L}M_{l}\times n\times K$, providing flexibility in controlling tunable parameters.
@@ 3 Experiments and Analysis
Table 1: Average accuracy@top1 over six image classification datasets using ResNet-34 in TIL setting (standard deviation over 3 seeds). RECAST surpasses non-adapter methods and enhances adapter-based approaches when combined. Parameters: $M$: model size, $R$: buffer, P: total parameters, $U$: unit attention mask, $G$: ghost modules The Learnt Parameters and Storage metrics are reported in a Per Task manner.
Method Learnt Params Storage Acc@top1(%)
ResNet-34 (He et al., 2015) 0.0 $M$ 65.2
@@
Baselines
To evaluate our proposed architecture, we have compared the performance against $15$ baselines in total, representing the aforementioned IL strategy categories. Among the popular We evaluate our architecture against $16$ baselines across multiple IL categories. From Regularization approaches, we include EWC (Kirkpatrick et al., 2017), which penalizes changes to critical network parameters, and LwF (Li & Hoiem, 2016), which uses previous model outputs as soft labels. For Replay Methods, we implement GDumb (Prabhu et al., 2020) with a 4400-sample memory buffer, and ECR-ACE Caccia et al. (2021) and DER++ (Buzzega et al., 2020), both using 120 samples per class. Reconfiguration methods are our primary focus, particularly adapter-based schemes that learn extra task-specific components. Among them, PiggyBack Mallya & Lazebnik (2018) uses binary masking schemes within a fixed architecture, while HAT Serra et al. (2018) learns soft attention mask. CLR Ge et al. (2023) is a purely adapter-based method, that introduces depthwise separable Ghost modules (Han et al., 2020) after each convolutional layer. We evaluate RECAST-ViT against LoRA-based approaches, that apply LoRA to various ViT components (Attention (Zhu et al., 2024) or Fully-connected layers (Chen et al., 2022)). We also compare with InfLoRA (Liang & Li, 2024), a hybrid between regularization, RoSA (Nikdan et al., 2024), which combines LoRA with sparse adaptation, DoRA (Liu et al., 2024b) that decomposes pretrained weight into magnitude and direction matrices, and a prompt-based adapter technique: L2P Wang et al. (2022). Finally, we include finetuned classifier layers as naive baselines.
Dataset
Following  Aljundi et al. (2019); Ge et al. (2023) we employ six diverse benchmarking datasets covering fine-grained to coarse image classification tasks across various domain (see Table 3 in Appendix A.1) including flowers (Nilsback & Zisserman, 2008), scenes (Quattoni & Torralba, 2009), birds (Wah et al., 2011), animals (Krizhevsky, 2009), vehicles (Maji et al., 2013), and other man-made objects (Krizhevsky, 2009). These datasets capture the diversity and inter-task variance essential for evaluating IL tasks. Images are standardized to $224\times 224$ pixels, applying basic transformations and dataset-specific normalization. The datasets are then pickled to ensure consistent comparisons across various experiments. More details are in Appendix A.1.
Table 2: Average accuracy@top1 over six image classification datasets using ViT-Small backbone in TIL setting (standard deviation over 3 seeds). RECAST improves baseline performance by  3% and enhances existing methods by$>1\%$. $M$: model parameters, $P$: total parameters, $r$: LoRA rank, $Pl$: prompt pool size, $To$: tokens per prompt, $D$: embedding dimension, $SpA$ : Sparse Adaptation. Learnt Parameters and Storage metric are reported in a Per Task manner (except L2P and InfLoRA)
Method Learnt Params Storage Acc@top1(%)
@@
RoSA (Nikdan et al., 2024)(SpA MLP) $6\times 10^{-6}P$ $M$ 84.5
RECAST (ours) + RoSA(SpA MLP) $6\times 10^{-6}P$ $M$ 85.4±$\pm{0.05}$
@@
Implementation Details
[…]
Figure 3: Performance comparison on Small ViT backbone: RECAST operates effectively in extremely few parameters range (24-96) where LoRA doesn’t function fruitfully. It also enhances LoRA’s performance when combined across all parameter ranges providing complementary benefits.
@@
Figure 4: Best accuracy@top1 averaged over six diverse image classification datasets, reported for ViT models of different scales. Performances are shown for Baseline Small-ViT fine-tuning, RECAST adapted ViT fine-tuning and RECAST adapted ViT finetuning with Coefficient finetuning
3.1 Results
RECAST demonstrates substantial improvements in continual learning across diverse neural network architectures. We evaluate its performance on ResNet-34 (Table 1) and ViT-Small (Table 2). Task-wise classification results are presented in Appendix Table 4, with ImageNet1k results across model scales and reconstruction schemes in Appendix 6. The results highlight RECAST’s ability to maintain high task accuracy while enhancing the performance of existing IL methods. Below, we compare RECAST with both non-adapter and adapter-based methods, focusing on parameter efficiency and storage requirements.
Table 1 showcases the performance of different methods using a ResNet-34 backbone. Traditional IL methods such as EWC (Lee et al., 2019) and LWF (Li & Hoiem, 2016) face drastic degradation as the number of tasks increases, demonstrating significant catastrophic forgetting with accuracies below $25\%$. GDUMB (Prabhu et al., 2020), which utilizes a large memory buffer, performs marginally better at $34.8\%$. Interestingly, HAT (Serra et al., 2018), despite learning unit-wise masks over the entire network, struggles to achieve high accuracy ($40.0\%$). This is likely because HAT was designed to allocate a fixed amount of network capacity to each task, and our diverse task suite challenges its ability to budget sufficient resources per dataset. In contrast, RECAST-FT attains $66.2\%$, surpassing the baseline ResNet-34 (65.2%) and outperforming other non-adapter methods, while finetuning only $3\times 10^{-6}$ of the original parameters. Among adapter methods, Piggyback (Mallya & Lazebnik, 2018) and CLR (Ge et al., 2023) show strong performance ($82.9\%$ and $79.9\%$ respectively). However, RECAST combined with these methods pushes the performance even further. RECAST + Piggyback achieves $84.1\%$, and RECAST + CLR reaches $80.9\%$, demonstrating RECAST’s ability to consistently improve adapter-based approaches.
Figure 5: The designing process of reconstruction is crucial to the IL performance of models. The plot above provides concrete evidence that Neural Mimicry with SmoothL1 is capable of generating backbones with most balanced template diversity and flexibility for Computer Vision Tasks
(a) Plot showing average classification accuracy of models reconstructed in different ways. VAE models performed slightly worse than the rest.
(b) Plot showing Diversity and adaptability measurements of models reconstructed in different ways. VAE models have shallow template diversity, and their internal information content is also less than the rest
@@ 3.1 Results
Table 2 summarizes results for methods using the ViT-Small (Dosovitskiy et al., 2020) backbone, highlighting RECAST’s strong performance. The baseline achieves $82.4\%$ accuracy, while RECAST-FT improves it to $85.2\%$—a $2.8$ percentage point gain with only $2\times 10^{-6}$ additional parameters. Adapter and reparameterization methods show strong performance with the ViT-Small backbone. InfLoRA (Liang & Li, 2024), MeLo (Zhu et al., 2024), AdaptFormer (Chen et al., 2022), RoSA (Nikdan et al., 2024), and DoRA (Liu et al., 2024b) all exceed $88\%$ accuracy. However, combining RECAST with these methods once again yields performance improvements. Combining RECAST with these enhances performance further, with RECAST + MeLo achieving $89.6\%$ and RECAST + AdaptFormer reaching $90.1\%$. In contrast, prompting methods like L2P (Wang et al., 2022), which rely on instance-level prompts, struggle in diverse domains, yielding a low accuracy of $35.9\%$.
Comparison of Resource Efficiency
@@
Tables 1 and 2 compare various incremental learning methods for ResNet-34 and ViT-Small backbones, respectively. We see that, baseline models (ResNet-34, ViT-Small) and parameter-free methods (EWC, LWF) maintain minimal storage (only the backbone model parameters) but suffer from performance degradation as task diversity increases. Rehearsal-based methods (GDUMB, DER++ & ER-ACE) and HAT offer improved performance at the cost of increased storage requirements, which may limit scalability for long-term learning. Recent parameter-efficient methods balance performance and resource usage. CLR learns $0.3\%$ of backbone parameters, requiring backbone storage plus task-specific depthwise modules. InfLoRA trains the full model for the first dataset and only $0.202\%$ for subsequent tasks. Other LoRA variants train less than $1\%$ per rank. While masking requires training a large number of parameters, both LoRA updates and binary masks (PiggyBack, RoSA Sparse updates) can be mathematically merged to prevent extra storage and communication costs during deployment. Our proposed RECAST method demonstrates unprecedented parameter efficiency, learning only $0.0002\%$ of the total backbone parameters while maintaining only the backbone model in storage. When combined with other efficient methods like MeLo or Adaptformer, RECAST achieves state-of-the-art performance (90.1% Acc@top1 for ViT-Small) with the base storage requirement. This reveals that RECAST not only enhances existing adapter methods but does so in ways that are complementary to their strengths. The impact is delineated clearly in Figure 3, in a range of taskwise parameter counts. It compares three adaptation approaches using ViT-Small backbone (82.4% baseline). Traditional LoRA adaptation (blue) becomes effective only above 464 parameters/task. RECAST (orange) achieves consistent 85% accuracy with just 24-96 parameters, showing marginal improvement beyond this range. The combined RECAST+LoRA (red) peaks at 91% accuracy, outperforming across all parameter ranges, including areas where LoRA alone falters. To scale traditional LoRA implementations to a negligibly small ($<150$) parameters, we utilized three techniques: (1) lowering LoRA rank, (2) sharing LoRA matrices across layers, and (3) binary masking for pruning. The figure reveals that RECAST can be used effectively in parameter ranges where other reparameterization methods might struggle.
@@ 3.2 Model Analysis
We analyze the effect of hyperparameters $G$, $n$, and $K$ (Subsection 2.3) through two experiments. First, we vary $n$ while keeping $K=1$, adjusting $G$ to maintain constant parameters ($G\propto\frac{1}{n}$), comparing three sharing schemes: Low (two-layer sharing), Balanced (four-layer sharing), and Extreme (all-layer sharing). Results are shown in Figure 7(a). Second, we fix $G$ and $n$ while varying $K$ (since $|C|\propto K$) as depicted in Figure 7(b). Both experiments use RECAST on MLP and Attention layers of a small ViT model, with marker sizes indicating relative coefficient sizes. Results reveal that adjusting template numbers and sharing schemes has a greater impact than linearly increasing $K$, emphasizing the importance of template diversity and specificity for task adaptation. Additional linear combinations show diminishing returns once sufficient coefficients are available (see the orange line in Figure 3).
Where do we apply RECAST?
@@
RECAST is applicable to various layers or modules (Appendix A.4.1), with notable insights highlighted in Figures 7(a) and 7(b). These figures illustrate the differing performance boosts when applying the method to attention layers versus MLP layers. MLP layers, though simpler, have more parameters that may introduce redundancy, making them more amenable to sharing. In contrast, attention layers are complex, involving multiple computational steps to model relationships between input parts rather than directly transforming representations. This complexity leads to a challenging optimization space, hindering effective template sharing and fine-tuning-based reparameterization. Empirical evidence in Figure 3 supports this: LoRA implementations for data points $1100$ and $13,000$ from AdaptFormer (Chen et al., 2022) show a performance spike when applied to MLP layers, indicating their greater impact compared to attention layers with fewer parameters (Zhu et al., 2024).
@@
(a) Increasing the number of templates, i.e: increasing the number of layers in a group, - both increases the number of coefficients and the intensity of template sharing across the architecture. In the plot, for Low sharing we set, $G=6$, for Balanced we set $G=3$ and for $Extreme$, $G=1$(one template bank is shared by all the layers of the network) This setting provides substantial performance boost.
(b) The number of coefficients can also change by increasing the number of coefficient sets, $K$. However, the gain in this case is not as pronounced - indicating that the model performance largely depends on the diversity of the templates.
@@
Figure 7: Plots comparing the strategies of changing coefficients. The size of the markers corresponds to the relative number of coefficients. The impact of increasing coefficients by varying the number of coefficient sets $K$, is much less prominent than changing it through grouping schemes.
What is the best way to reconstruct weights?
@@
Section 2.2 described the weight reconstruction methodology for RECAST. Earlier we noted that model expressivity depends more on template bank diversity. To investigate this, we generated five models with the same configurations (6 groups, 2 layers each), varying the reconstruction objective: 1) SmoothL1 loss, 2) MSE loss, 3) SmoothL1 loss with coefficient noise, 4) VAE with MSE loss, and 5) VAE with SmoothL1 loss. All models achieved at least $98.5\%$ reconstruction similarity (Table 6).Figure 6(a) shows that SmoothL1 loss (with or without perturbation) provides the most benefit during reparameterization & VAE models provide the least. To understand this, we calculated template diversity within groups using Frobenius Norm (diversity metric) and Entropy (versatility metric) (See Appendix A.3) Just like their classification performance, SmoothL1 loss shows roughly similar internal patterns with or without noise. The high Frobenius norm for Neural Mimicry (Algo. 2) with SmoothL1 loss indicates greater template diversity than the rest of the approaches. MSE models showed slightly lower Frobenius norms but higher entropy, suggesting templates are versatile and not overly dependent on dominant components. In contrast, VAEs showed shallow template diversity, limiting adaptability, and lower entropy, indicating reduced versatility. Figure 9 (Appendix A.3) illustrates higher intra-group coefficient similarity in VAE-based models, suggesting multiple layers perform similar operations.

### v2

@@ 3.3 Parametric Scaling Strategy
RECAST’s design is governed by three parameters: the number of groups ($G$), templates per group ($n$), and coefficient sets ($K$), balancing model expressiveness and efficiency. Increasing $G$ improves specialization, while increasing $n$ increases expressiveness within groups. The parameter-sharing scheme allows $K$ to expand weight configurations without significantly increasing parameters. This design achieves substantial memory savings, $\mathbb{S}=L\times M_{l}\times d^{2}-(G\times n\times d^{2}+L\times M_{l}\times n\times K)$. Here, $M_{l}$ refers to the number of target modules per layer, and we can approximate the savings as: $\mathbb{S}\approx L\times M_{l}\times d^{2}-G\times n\times d^{2}=d^{2}(L\times M_{l}-G\times n)$. The savings primarily stem from replacing $L\times M_{l}$ weight matrices with $G\times n$ templates. In our experiments $G$ and $n$ are calibrated to match the baseline model’s parameter count but can be adjusted for desired compression. Additionally, multiple coefficient sets enable a combinatorially large number of weight matrices ($n^{K}$) per module, covering extensive parameter spaces even with small $n$. The total learnable coefficients are $\sum_{l=1}^{L}M_{l}\times n\times K$, providing flexibility in controlling tunable parameters.
@@ 4 Experiments and Analysis
Baselines. To evaluate our proposed architecture, we have compared the performance against $16$ baselines in total, representing the aforementioned categories of IL strategies. From Regularization approaches, we include EWC (Kirkpatrick et al., 2017), which penalizes changes to critical network parameters, and LwF (Li & Hoiem, 2016), which uses previous model outputs as soft labels. For Replay Methods, we implement GDumb (Prabhu et al., 2020) with a 4400-sample memory buffer, and ECR-ACE (Caccia et al., 2022) and DER++ (Buzzega et al., 2020), both using 120 samples per class. Reconfiguration methods are our primary focus as they are most similar to our approach, particularly adapter-based schemes that learn extra task-specific components. Among them, PiggyBack (Mallya et al., 2018) uses binary masking schemes within a fixed architecture, while HAT Serra et al. (2018) learns a soft attention mask. CLR Ge et al. (2023) is a purely adapter-based method, that introduces depthwise separable Ghost modules (Han et al., 2020) after each convolutional layer. We also consider several LoRA-based approaches that apply LoRA to various ViT components (Attention (Zhu et al., 2024) or Fully-connected layers (Chen et al., 2022)). In addition, we compare with InfLoRA (Liang & Li, 2024), which has regularization and reconfiguration components, RoSA (Nikdan et al., 2024), which combines LoRA with sparse adaptation, DoRA (Liu et al., 2024b) that decomposes pretrained weight into magnitude and direction matrices, and a prompt-based adapter technique: L2P (Wang et al., 2022b). Finally, we include finetuned classifier layers as naive baselines.
Datasets. Following standard experiment setups in similar works by  Aljundi et al. (2019); Ge et al. (2023) we employ six diverse benchmarking datasets covering fine-grained to coarse image classification tasks across various domain including flowers (Nilsback & Zisserman, 2008), scenes (Quattoni & Torralba, 2009), birds (Wah et al., 2011), animals (Krizhevsky & Hinton, 2009), vehicles (Maji et al., 2013), and other man-made objects (Krizhevsky & Hinton, 2009). These datasets capture the diversity and inter-task variance essential for evaluating IL tasks. Images are standardized to $224\times 224$ pixels, applying basic transformations and dataset-specific normalization. Additional details are in Appendix A.1 Table 3 .
Table 1: TIL accuracy averaged over six datasets using a ResNet-34 comparing non-adapter (top) and adapter methods (bottom). RECAST surpasses non-adapter methods and enhances adapter-based approaches when combined. Parameters: $M$: model size, $R$: buffer, P: total parameters, $U$: unit attention mask, $G$: ghost modules Task Parameters and Storage metrics are reported Per Task
Method Task Params Storage Acc@top1(%)
ResNet-34 (He et al., 2016) 0.0 $M$ 65.2
@@
Metrics. We report Top-1 accuracy averaged over our six datasets. For RECAST, we report the mean and standard deviation averaged over 3 runs. See Appendix Table 4 for per-task performance.
[…]
4.1 Results
Table 1 showcases the performance of different methods using a ResNet-34 backbone. Traditional IL methods such as EWC (Lee et al., 2019) and LWF (Li & Hoiem, 2016) show significant catastrophic forgetting, with performance drops of approximately $66\%$ and $70\%$ respectively, compared to the baseline ResNet-34. GDUMB (Prabhu et al., 2020), ER-ACE (Caccia et al., 2022), despite utilizing a large memory buffer, only perform slightly better than these traditional methods but still lag significantly behind simple finetuning. HAT (Serra et al., 2018), which learns unit-wise masks over the entire network, achieves further improvement over GDUMB but still lags $39\%$ behind the baseline. This underperformance is likely because HAT was designed to allocate fixed network capacity to each task, struggling with our diverse task suite. In contrast, RECAST improves upon the baseline ResNet-34 by $\sim 2\%$, while fine-tuning only 0.0003% of the original parameters. This represents a significant efficiency advantage while still delivering performance gains. Among adapter methods, Piggyback (Mallya et al., 2018) and CLR (Ge et al., 2023) show strong performance, exceeding the baseline by $27\%$ and $23\%$ respectively. Combined with these methods, RECAST further enhances their effectiveness. RECAST + Piggyback achieves $\sim 1.5\%$ improvement over Piggyback alone. Similarly, RECAST + CLR also shows a $\sim 1.5\%$ gain over standalone CLR. These consistent improvements across different adapter-based approaches demonstrate the versatility and effectiveness of RECAST as a complementary technique.
Table 2: TIL accuracy averaged over six datasets using a ViT-Small comparing non-adapter (top) and adapter methods (bottom). RECAST improves baseline accuracy by 3 points and enhances existing methods $>1\%$. $M$: model parameters, $P$: total parameters, $r$: LoRA rank, $Pl$: prompt pool size, $To$: tokens per prompt, $D$: embedding dimension, $SpA$ : Sparse Adaptation. Task Parameters and Storage metric are reported in a Per Task manner (except L2P and InfLoRA)
Method Task Params Storage Acc@top1(%)
@@ 4.1 Results
RECAST (ours) + RoSA(SpA MLP) $6\times 10^{-6}P$ $M$ 85.4±$\pm{0.05}$
@@
Figure 3: Averaged best Top-1 accuracy across six datasets for ViT models of varying scales, comparing Baseline, RECAST, and RECAST with Coefficient fine-tuning (FT) approaches.
Figure 4: Comparing averaged best Top-1 accuracy across six datasets for ViT-Small across various model configurations. We find RECAST excels in ultra-low parameter ranges (24-96) where LoRA struggles. RECAST w/ LoRA improves performance across all parameter ranges, offering complementary advantages.
@@
Table 2 summarizes the results for methods using the ViT-Small (Dosovitskiy et al., 2020) backbone, further emphasizing RECAST’s strong performance. RECAST improves the baseline ViT-Small by ${}\sim 3.5\%$, adding only $0.0002\%$ task parameters. This represents a significant enhancement with minimal computational overhead. Adapter and reparameterization methods all show strong performance improvements over the baseline: RoSA (Nikdan et al., 2024) (with SpA MLP) improves by $2.5\%$, while InfLoRA (Liang & Li, 2024), MeLo (Zhu et al., 2024), AdaptFormer (Chen et al., 2022), DoRA (Liu et al., 2024b), and RoSA (Full - Att.) all deliver improvements in the $7.5-8.5\%$ range. However, combining RECAST with these methods yields additional gains. Both RECAST + RoSA (SpA MLP) and RECAST + AdaptFormer, show a $\sim 1.5\%$ gain over standalone RoSA and AdaptFormer alone, and RECAST + MeLo achieves a $1.0\%$ improvement over MeLo. In contrast, prompting methods like L2P (Wang et al., 2022b), which rely on instance-level prompts, underperform the baseline by $56\%$, struggling to handle diverse domains effectively.
Figure 4.1 reports the performance of using RECAST over a range of architecture sizes including ViT-Small, ViT-Base, and ViT-Large, demonstrating that our approach generalizes. In addition, we also show that our approach generalizes across pretraining datasets by demonstrating our approach an boost performance using a DINO backbone Caron et al. (2021).
Figure 5: Plot showing average classification accuracy of models reconstructed in different ways. VAE-reconstructed models performed slightly worse than the rest, with Smooth L1 loss providing the best performance
Figure 6: Plot showing Diversity (Frobenius Norm) and adaptability measurements (Entropy) of models reconstructed in different ways. VAE models have shallow template diversity and low internal information content
@@
Comparison of Resource Efficiency. Tables 1 and 2 compare various incremental learning methods for ResNet-34 and ViT-Small backbones, respectively. We see that parameter-free methods (EWC, LWF) maintain minimal storage (only the backbone model parameters) but suffer from performance degradation as task diversity increases. Rehearsal-based methods (GDUMB, DER++ & ER-ACE) and HAT offer improved performance at the cost of increased storage requirements, which may limit scalability for long-term learning. Recent parameter-efficient methods balance performance and resource usage. CLR learns $0.3\%$ of backbone parameters, requiring backbone storage plus task-specific depthwise modules. InfLoRA trains the full model for the first dataset and only $0.2\%$ for subsequent tasks. Other LoRA variants train less than $1\%$ per rank. While masking requires training a large number of parameters, both LoRA updates and binary masks (PiggyBack, RoSA Sparse updates) can be mathematically merged to prevent extra storage and communication costs during deployment. Our proposed RECAST method demonstrates unprecedented parameter efficiency, learning only $0.0002\%$ of the total backbone parameters while maintaining only the backbone model in storage. When combined with other efficient methods like MeLo or Adaptformer, RECAST achieves state-of-the-art performance (90.1% Acc@top1 for ViT-Small) with the base storage requirement. This reveals that RECAST not only enhances existing adapter methods but does so in ways that are complementary to their strengths. The impact is delineated clearly in Figure 4, in a range of taskwise parameter counts. It compares three adaptation approaches using ViT-Small backbone (82.4% baseline). Traditional LoRA adaptation (blue) becomes effective only above 464 parameters/task. RECAST (orange) achieves consistent 85% accuracy with just 24-96 parameters, showing marginal improvement beyond this range. The combined RECAST+LoRA (red) peaks at 91% accuracy, outperforming across all parameter ranges, including areas where LoRA alone falters. To scale traditional LoRA implementations to a negligibly small ($<150$) parameters, we utilized three techniques: (1) lowering LoRA rank, (2) sharing LoRA matrices across layers, and (3) binary masking for pruning. The figure reveals that RECAST can be used effectively in parameter ranges where other reparameterization methods might struggle.
@@ 4.2 Model Analysis
How does the number of coefficients impact performance? We analyze the effect of hyperparameters $G$, $n$, and $K$ (Subsection 3.3) through two experiments. First, we vary $n$ while keeping $K=1$, adjusting $G$ to maintain constant parameters ($G\propto\frac{1}{n}$), comparing three sharing schemes: Low (two-layer sharing), Balanced (four-layer sharing), and Extreme (all-layer sharing). Results are shown in Figure 8. Second, we fix $G$ and $n$ while varying $K$ (since $|C|\propto K$) as depicted in Figure 8. Both experiments use RECAST on MLP and Attention layers of a small ViT model, with marker sizes indicating relative coefficient sizes. Results reveal that adjusting template numbers and sharing schemes has a greater impact than linearly increasing $K$, emphasizing the importance of template diversity and specificity for task adaptation. Additional linear combinations show diminishing returns once sufficient coefficients are available (see the orange line in Figure 4).
@@
Where do we apply RECAST? RECAST is applicable to various layers or modules (Appendix A.4.1), with notable insights highlighted in Figures 8 and 8. These figures illustrate the differing performance boosts when applying the method to attention layers versus MLP layers. MLP layers, though simpler, have more parameters that may introduce redundancy, making them more amenable to sharing. In contrast, attention layers are complex, involving multiple computational steps to model relationships between input parts rather than directly transforming representations. This complexity leads to a challenging optimization space, hindering effective template sharing and reparameterization. Figure 4 confirms this pattern, showing performance spikes when smaller LoRA matrices target MLP layers (Chen et al., 2022) rather than attention mechanisms (Zhu et al., 2024).
@@
Figure 7: Increasing template numbers (layers in a group) enhances coefficient count and template sharing across the architecture. Low sharing uses G=6, Balanced G=3, and Extreme G=1 (one template bank shared by all network layers). The last setting delivers significant performance improvements.
@@
Figure 8: Comparing the strategies of changing coefficients. The size of the markers corresponds to the relative number of coefficients. The impact of increasing coefficients by varying the number of coefficient sets $K$, is much less prominent than changing it through grouping schemes.
@@
What is the best way to reconstruct weights? Section 3.2 described the weight reconstruction methodology for RECAST. Earlier we noted that model expressivity depends more on template bank diversity. To investigate this, we generated five models with the same configurations (6 groups, 2 layers each), varying the reconstruction objective: 1) SmoothL1 loss, 2) MSE loss, 3) SmoothL1 loss with coefficient noise, 4) VAE with MSE loss, and 5) VAE with SmoothL1 loss. All models achieved at least $98.5\%$ reconstruction similarity (Appendix Table 6). Figure 5 shows that SmoothL1 loss (with or without perturbation) provides the most benefit during reparameterization & VAE models provide the least. To understand this, we calculated template diversity within groups using Frobenius Norm (diversity metric) and Entropy (versatility metric) (See Appendix A.3) Just like their classification performance, SmoothL1 loss shows roughly similar internal patterns with or without noise. The high Frobenius norm for Neural Mimicry (Algorithm 2) with SmoothL1 loss indicates greater template diversity than the rest of the approaches. MSE models showed slightly lower Frobenius norms but higher entropy, suggesting templates are versatile and not overly dependent on dominant components. In contrast, VAEs showed shallow template diversity, limiting adaptability, and lower entropy, indicating reduced versatility. Figure 10 (Appendix A.3) illustrates higher intra-group coefficient similarity in VAE-based models, suggesting multiple layers perform similar operations.